1. Park officials make predictions of times to the next eruptiom of a particular geyser, and collect data for the errora (minutes) in those predictions. The display from technology available below results from using the prediction errors to test the claim that the mean prediction error is equal to zero. Comment on the accuracy of the predictions. Use a 0.05 significance level. Identify the null and alternative hypothesis, test statistic, and P-Value.

2. In a test of the effectiveness kf harlicnfor lowering cholesterol, 81 subjects were treated with raw garlic. Cholesterol levels were measured before and after the treatment. The changes in their levels of LDL cholesterol have a mean of 0.3 and a standard deviation of 2.08. Use a 0.01 significance level to test the claim that with garlic treatment, the mean change is greater than 0. Assume a simple random test has been selected. Identify the null and alternative hypothesis, test statistic and P-value.

3. A data set lists earthquake depths. The summary statistics are n=400, x=4.45 km, s= 4.87 km. Use a 0.01 significance level to test the claim of a seismologist that these earthquakes are from a population with a mean equal to 4.00. Assume that a simple randomnwample has been selected. Identify the null and alternatice hypothesis, test statistic, and P-Value.
Answer #1

Answer:

3.

Given,

sample n = 400

mean x = 4.45

standard deviation = 4.87

Ho : u = 4

Ha : u != 4

Here it is two tailed test

test consider,

t = (x - u)/(s/sqrt(n))

substitute values

= (4.45 - 4)/(4.87/sqrt(400))

t = 1.85

degree of freedom = n - 1

= 400 - 1

= 399

P value = 0.065053

Here we observe that, p value > 0.01, so we fail to reject Ho.

So there is no enough evidence to support the claim that of a seismologist that these earthquakes.

Please post the remaining questions as separate post. Thank you.